# Review: Quantum Metrology and Sensing with Many-Body Systems

### v1

@@ VIII.1 Dynamical quantum phase transitions quantified by Loschmidt echo
It is discovered that if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belongs to two-different equilibrium phases, then the ${\cal G}(t)$ can be non-analytic at certain time $t{=}t_{c}$. On the other hands, if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belongs to the same equilibrium phase then ${\cal G}(t)$ is a regular function of time Heyl (2018). There are a few peculiarities to these cases Jafari and Johannesson (2017) as well as extension in non-Hermitian systems Zhou et al. (2018); Jing et al. (2024). Such dynamical phase transitions has been proposed to provide quantum enhancement in parameter estimation González et al. (2022). The authors have proposed a negatively charged Nitrogen-Vacancy (NV-) center in a nearby Carbon-13 nuclear spins as a platform for studying dynamical phase transitions. First, they have shown that the system undergoes a dynamical phase transition by a suitable choice of Gaussian time-varying magnetic field. Then it is shown that the Fisher information, for estimating coupling strength between the two Carbon-13 nuclear spins, improves at $t{=}t_{c}$ as compared to the regular times.
@@ IX Non-Equilibrium Quantum Sensing: Discrete Time Crystal Phase Transitions
Apart from dynamical phase transitions, discussed in Sec. VIII, there are other categories of phase transitions which occur through non-equilibrium dynamics of many-body systems. An important class of such systems are discrete-time crystals (DTCs) which are defined in an isolated non-equilibrium quantum many-body system that undergoes periodic drives. Breaking the time transnational symmetry results in the emergence of a new phase of matter known as the DTC phase with indefinite persistent oscillations Sacha (2015); Khemani et al. (2016); Else et al. (2016). Recently the long-range spatial- and time-ordered dynamics in the DTC phase have been harnessed in measuring an AC filed Iemini et al. (2024b); Gribben et al. (2024). The optimal performance for long periods along with intrinsic robustness to the imperfections in the sensing protocol based on DTC phase allows one to hit the quantum standard limit in estimating the AC field. However, obtaining a true sensing enhancement based on a resource such as system size requires a wised mechanism for establishing DTC. In Ref. Yousefjani et al. (2024b) the proposed mechanism results in a stable DTC with period-doubling oscillations that persist indefinitely even in finite-size systems. The time-dependent Hamiltonian that governs the dynamic in the considered spin-$1/2$ chain is
@@ X Non-Equilibrium Quantum Sensing: Floquet Phase Transition
Quantum metrologically advantageous states can also be generated using the application of external time-periodic driving to the system. In this case, the frequency of the external driving can resonate with the lowest energies of the unperturbed systems. This resonance can be harnessed for quantum enhanced sensing Lang et al. (2015); Wang et al. (2022b); Jiang et al. (2022); Pang and Jordan (2017); Zhou et al. (2020); Mishra and Bayat (2021, 2022); Yu et al. (2022); Ivanov (2021). This resonance phenomena has also been observed in gaped many-body quantum systems when the frequency of external driving matches with the lowest energy gap of the unperturbed Hamiltonian. Russomanno et al. (2016); Mishra and Bayat (2021, 2022). Various important physical features of systems with time-periodic influence can be extracted by employing Floquet formalism. For a time-periodic Hamiltonian, $H(t+\tau)=H(t)$, with periodicity $\tau=2\pi/\omega$, the solution of the Schrödinger equation follows from the Floquet theorem. The Floquet theorem gives an ansatz of the form $|\Psi(t)\rangle=\sum_{\ell}e^{-i\mu_{\ell}t}|\Phi_{\ell}(t)\rangle$. Here $\mu_{\ell}^{\prime}s$ are eigenphase (the Floquet phase), $|\Phi_{\ell}(t)\rangle$ the Floquet modes, and $\ell\in\mathbb{Z}$. Substituting the ansatz to the Schrödinger equation, $H(t)|\Psi(t)\rangle=-i\hbar\frac{\partial|\Psi(t)\rangle}{\partial t}$, the Floquet modes satisfy
@@
Properties of Floquet modes $\{|\Phi_{\ell}(t)\}$ and Floquet phases $\{\mu_{\ell}\}$ have been explored for sensing periodic magnetic fields due to clusters of few qubits Lang et al. (2015). Here, the effective magnetic field due to the cluster of spins is detected using a single qubit as a sensor. The initial state of the system is $|\psi(0)\rangle{=}\frac{1}{\sqrt{2}}(\ket{\uparrow}{+}\ket{\downarrow})\otimes|\mathcal{B}(0)\rangle$, where |$\ket{\mathcal{B}(0)}$ is the initial state of the detected spin cluster at time $t=0$ and $\ket{\uparrow}$ , $\ket{\downarrow}$ are the computational basis states of the electronic spin, respectively. At later times, due to the interaction with the sensor (the qubit state), the spin cluster state becomes entangled with the sensor. The magnetic field signal is detected from the data of temporal coherence of the time-evolved state of the detected spin cluster. The coherence can be expressed in terms of the eigenvalues and eigenstates of the one-period unitary evolution operator or Floquet operator. The coherence displays a dip which occurs at the avoided crossings of the Floquet eigenstates where $e^{i\mu_{\ell}}\approx e^{-i\mu_{k}}$ or when the Floquet gap, $\Delta_{F}=\mu_{\ell}-\mu_{k}$, vanishes. Thus, by measuring the location of the coherence dip, the energy-difference between Floquet states can be observed with precision higher than the one predicted by the average Hamiltonian theory. However, this method can only be applicable for a sensing around the resonance frequencies. Sensing of arbitrary frequency fields has been proposed using a single qubit as a sensor in Floquet systems Wang et al. (2022b). In addition, periodically driven systems and the transition of Floquet eigenphases have been used for magnetic-field signal amplification Jiang et al. (2022). In Ref. Pang and Jordan (2017), an optimal control method is developed to estimate the amplitude, $B$, of a time-periodic magnetic field $B(t)=B(\cos(\omega t)\sigma_{x}+\sin(\omega t)\sigma_{z})$. It is shown here that the QFI with respect to $B$ can achieve super-Heisenberg scaling with interrogation time, $t$, by applying an optimal control pulse as compared to the one without control. A solid state sensor with an ensemble of high density Nitrogen Vacancy (NV) centers in diamonds has been investigated both theoretically and experimentally by Floquet formalism  Zhou et al. (2020). The system, here, is a dense ensemble of electronic spins. Each electronic spin act as an independent probe for the external signal. Thus, one would expect that, by increasing the number of particles in the system, the sensitivity can be improved. The improvement, however, is severely hindered by interaction between the spins, since a large spins are confined in a small volume, the spin-spin interaction is unavoidable. By applying a series of dynamical decoupling pulses, the interaction between the spins has been suppressed, which results in a larger coherence time. In this way the standard quantum limit is shown to be restored which is lost due to spin-spin interaction.
@@ XI.1 Profiting from dissipation
Since non-classical states are crucial for countless applications and theoretical advances, it would be highly desirable to have a mechanism where instead of avoiding dissipation, one could harness the very process of dissipation to one’s advantage. The seminal work by Alfred Kastler in 1950 pointed in that direction Kastler, Alfred (1950). The process, known as optical pumping, prepares any initial quantum state into a well-defined target state via selection rules of spontaneous emission. Building on Kastler’s work, coherent population trapping was experimentally demonstrated in the 1970s by Arimondo and Orriols Alzetta et al. (1976); Arimondo and Orriols (1976). In a nutshell, a three-level energy configuration with transitions levels $|a\rangle\longleftrightarrow|b\rangle$ and $|a\rangle\longleftrightarrow|c\rangle$ and Hamiltonian $H=\hbar\Omega(|a\rangle\langle b|-|a\rangle\langle c|+h.c)$ allows for a quantum state with a zero eigenvalue (a dark state). Remarkably, such a dark state, denoted as $|\psi_{d}\rangle\sim|b\rangle+|c\rangle$, decouples from the external pumping field $\Omega$, thus allowing excitations that reach that energy level to remain there. In contrast, the bright state, denoted as $|\psi_{b}\rangle\sim|b\rangle-|c\rangle$, oscillates with the driving field $\Omega$ and the state $|a\rangle$. Indeed, this idea lies at the heart of laser cooling Aspect et al. (1988) as the ground state represents another specific quantum state. The paradigm shift of harnessing dissipation as a resource rather than a drawback has led to systematic studies of decoherence itself Myatt et al. (2000), the ability to prepare a quantum state from any arbitrary initial state Poyatos et al. (1996), the use of atomic engineered reservoirs to create two-mode Einstein-Podolsky-Rosen-entangled states Pielawa et al. (2007, 2010), protection against noise through the engineering of pointer states Carvalho et al. (2001), dissipative quantum computation Verstraete et al. (2009), the creation of many-body states and non-equilibrium quantum phases Diehl et al. (2008), quantum memories using dissipative mechanisms Pastawski et al. (2011), the generation of Bell states on an open-system quantum simulator Barreiro et al. (2011), and the possibility to enhance sensing precision nearly to the Heisenberg limit Goldstein et al. (2011).
@@ XI.3 Dissipative phase transition as a sensing resource
Intense theoretical work has been pursued to understand the critical behavior at the dissipative phase transition Rota et al. (2017); Vicentini et al. (2018); Letscher et al. (2017); Overbeck et al. (2017); Capriotti et al. (2005); Eisert and Prosen (2010); Sieberer et al. (2014); Garbe et al. (2020); Ivanov (2020), including photonic systems Carmichael (2015); Benito et al. (2016); Mendoza-Arenas et al. (2016); Casteels et al. (2016); Bartolo et al. (2016); Casteels et al. (2017); Casteels and Ciuti (2017); Foss-Feig et al. (2017); Biondi et al. (2017); Biella et al. (2017); Savona (2017); Sánchez Muñoz et al. (2018), lossy polariton condensates Sieberer et al. (2013, 2014), and spin systems Kessler et al. (2012); Lee et al. (2013); Jin et al. (2016); Lee et al. (2011); Chan et al. (2015); Maghrebi and Gorshkov (2016); Overbeck et al. (2017). In addition, exact steady-state expressions Lawande et al. (1981); Puri et al. (1980); Bartolo et al. (2016) and approximate methods for their evaluation based on variational approaches have been pursued Weimer (2015); Raghunandan et al. (2018). The presence of phase symmetry breaking in a single qubit-laser system Fernández-Lorenzo and Porras (2017), semiclassical first-order dissipative phase transition in Kerr parametric oscillators with parity symmetry breaking Heugel et al. (2019), and second-order transitions in the two-photon Kerr resonator Di Candia et al. (2023). The universality class of driven-dissipative systems has also been extensively addressed Marcuzzi et al. (2014); Carollo et al. (2019b); Sieberer et al. (2013); Marino and Diehl (2016); Maghrebi and Gorshkov (2016) with in-depth analysis of the phase diagram and critical exponents of a dissipative transverse Ising spin chain Werner et al. (2005). From a geometric perspective, the Uhlmann curvature has been proposed to shed light upon the nature of criticality Carollo et al. (2018b), and a recent method based on the coherent anomaly approach has been proposed for determining critical exponents Jin et al. (2021).
@@
Quantum-enhanced sensitivity using the BTC probe of Eq. (76) can be evidenced in Fig. 15(a) via QFI analysis Montenegro et al. (2023). As seen from the figure, the maximum of the QFI $\mathcal{F}_{Q}^{\mathrm{max}}$ grows polynomially with the system size $N$. A fitting function of the form $\mathcal{F}_{Q}^{\mathrm{max}}=aN^{b}+c$ ($c\rightarrow 0$) reveals a coefficient $b=1.345$. Interestingly, in Fig. 15(a), a simple measurement basis of the total magnetization also scales super-linearly with the system size $N$. A fitting function of the same form as above exhibits quantum-enhanced sensitivity with a coefficient $b=1.338$. To show the performance between the QFI and the classical Fisher information, in Fig. 15(b), the ratio between their maximum values is plotted, showing an asymptotic performance of 90% —a fair fraction achieved with a sub-optimal measurement basis.
@@ XIII Non-Equilibrium Quantum Sensing: Sequential Measurements Metrology
Measurement represents another distinct feature that sets apart the classical from the quantum world. Indeed, the concept of wave-function collapse due to projective measurements on the quantum system has sparked several discussions within quantum theory itself Machida and Namiki (1980a, b); Furry (1936); Zeh (1970). In quantum many-body probes with partial accessibility, measurements can only be conducted locally on a subsystem. However, in these scenarios, even though the measurement is local, the wave function collapse is global and affects the entire system’s wave function. This intriguing phenomenon has been the subject of intensive studies Busch et al. (1990); Schmidt and Gemmer (2020); Ban (2021); Skinner et al. (2019b); Block et al. (2022a); Benoist et al. (2023); Haapasalo et al. (2016); Benoist et al. (2019); Burgarth et al. (2014); Rybár and Ziman (2015), such as a unique feature for quenching many-body systems Bayat (2017); Pouyandeh et al. (2014); Ma et al. (2018); Bayat et al. (2018), leading to potential new types of phase transitions Choi et al. (2020); Feng et al. (2023); Minato et al. (2022); Skinner et al. (2019a); Gullans and Huse (2020); Block et al. (2022b).
@@
where $\bm{\sigma}_{j}{=}(\sigma_{j}^{x}{,}\sigma_{j}^{y}{,}\sigma_{j}^{z})$ represents a vector composed of Pauli matrices acting on qubit site $j$, $J$ denotes the exchange interaction, and $\bm{B}{=}(B_{x}{,}0{,}B_{z})$ stands for the local magnetic field to be estimated located at first site. For the sake of simplicity, a ferromagnetic state $|\psi(0)\rangle=|\downarrow\downarrow\downarrow\ldots\rangle$ is probed at site $N$ in the $\sigma_{z}$ basis at regular time intervals $J\tau=N$. The latter allows enough time for the quantum correlations to spread throughout the spin chain. It has been shown Montenegro et al. (2022) that a synchronization-like magnetization dynamics occurs between the first and last sites. This opens up the possibility of extracting information about the local magnetic field (located at the first site) by observing the dynamics of the readout qubit (located at the last $N$ site). Remote sensing holds special significance in biology, where non-invasive sampling of organic tissues is desirable to protect the biological material (see quantum illumination techniques Albarelli et al. (2020)).
@@
where $\mathcal{F}_{\lambda}^{(j)}$ represents the classical Fisher information at step $j$, $\Delta\mathcal{F}_{\lambda}^{(n)}$ is the increment of the classical Fisher information after performing one more measurement following the recording of $(n{-}1)$ measurements, and $f_{\lambda}^{\boldsymbol{\gamma},(n)}$ is the classical Fisher information obtained from the $n$-th measurement $p(\gamma_{n})$ in trajectory $\boldsymbol{\gamma}$. Utilizing this approach, one can examine the behavior of the classical Fisher information for arbitrarily large numbers of sequential measurements. Conducting local projections on a quantum many-body probe leads to probe’s finite memory, attributed to repeated local projections and free evolution, resulting in a rank-1 matrix Yang et al. (2023). This probe’s finite memory is linked to the transition from super-linear to linear scaling of classical Fisher information. After losing memory of an early state, classical Fisher information scales linearly with $n_{\mathrm{seq}}$, limiting the extractable information capacity of the quantum probe in sequential measurements.
For the quantum many-body probe described in Eq. (82), Figs. 18(a)-(b) demonstrate a correspondence between the transition from super-linear to linear scaling of classical Fisher information and the probe’s finite memory. Specifically, the classical Fisher information increment $\Delta F_{B}^{(n_{\mathrm{seq}})}$ becomes nearly constant at around the same $n_{\mathrm{seq}}$ where the averaged fidelity $\langle F\rangle_{\mathrm{traj}}$ also becomes nearly constant. This averaged fidelity measures the similarity between two distinct initial states following the same trajectory Yang et al. (2023) (a single trajectory is composed of $n_{\mathrm{seq}}$ measurement steps). In Fig. 18(c), the classical Fisher information $\mathcal{F}_{B}^{(n_{\mathrm{seq}})}$ is plotted as a function of $n_{\mathrm{seq}}$. This plot clearly shows a transition from non-linear to linear behavior, in agreement with the quantum-enhanced sensing shown for short sequential measurements Montenegro et al. (2022) and the linear scaling with indirect methods for large $n_{\mathrm{seq}}$ Burgarth et al. (2014). The Monte Carlo-based methodology discussed above for arbitrarily large $n_{\mathrm{seq}}$ it has been addressed for both quantum many-body probes and light-matter probes Yang et al. (2023).
Figure 18: Sequential measurements metrology for the quantum-many body probe of Eq. (82) and large number of consecutive measurements $n_{\mathrm{seq}}$: $(a)$ Classical Fisher information increment $\Delta F_{B}^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$, $(b)$ Fidelity averaged over $10^{4}$ different trajectories (each trajectory evaluates the fidelity between two distinct initial states following the same trajectory) as a function of $n_{\mathrm{seq}}$, $(c)$ classical Fisher information $\mathcal{F}_{B}^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$. Figure taken from Yang et al. (2023).

### v2

@@ VIII.1 Dynamical quantum phase transitions quantified by Loschmidt echo
It is discovered that if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belongs to two-different equilibrium phases, then the ${\cal G}(t)$ can be non-analytic at certain time $t{=}t_{c}$. On the other hands, if $|\Psi(0)\rangle$ and $|\Psi(t)\rangle$ belongs to the same equilibrium phase then ${\cal G}(t)$ is a regular function of time Heyl (2018). There are a few peculiarities to these cases Jafari and Johannesson (2017) as well as extension in non-Hermitian systems Zhou et al. (2018); Jing et al. (2024). Such dynamical phase transitions have been experimentally realized in large quantum simulators (Zhang et al., 2017) and has also been proposed to provide quantum enhancement in parameter estimation González et al. (2022), where the authors have proposed a negatively charged Nitrogen-Vacancy (NV-) center in a nearby Carbon-13 nuclear spins as a platform for studying dynamical phase transitions. First, they have shown that the system undergoes a dynamical phase transition by a suitable choice of Gaussian time-varying magnetic field. Then it is shown that the Fisher information, for estimating coupling strength between the two Carbon-13 nuclear spins, improves at $t{=}t_{c}$ as compared to the regular times.
@@ IX Non-Equilibrium Quantum Sensing: Discrete Time Crystal Phase Transitions
Apart from dynamical phase transitions, discussed in Sec. VIII, there are other categories of phase transitions which occur through non-equilibrium dynamics of many-body systems. An important class of such systems are discrete-time crystals (DTCs) which are defined in an isolated non-equilibrium quantum many-body system that undergoes periodic drives. Breaking the time transnational symmetry results in the emergence of a new phase of matter known as the DTC phase with indefinite persistent oscillations Sacha (2015); Khemani et al. (2016); Else et al. (2016); Yoshinaga et al. (2022). Recently the long-range spatial- and time-ordered dynamics in the DTC phase have been harnessed in measuring an AC filed Iemini et al. (2024b); Gribben et al. (2024). The optimal performance for long periods along with intrinsic robustness to the imperfections in the sensing protocol based on DTC phase allows one to hit the quantum standard limit in estimating the AC field. However, obtaining a true sensing enhancement based on a resource such as system size requires a wised mechanism for establishing DTC. In Ref. Yousefjani et al. (2024b) the proposed mechanism results in a stable DTC with period-doubling oscillations that persist indefinitely even in finite-size systems. The time-dependent Hamiltonian that governs the dynamic in the considered spin-$1/2$ chain is
@@ X Non-Equilibrium Quantum Sensing: Floquet Phase Transition
Quantum metrologically advantageous states can also be generated using the application of external time-periodic driving to the system. In this case, the frequency of the external driving can resonate with the lowest energies of the unperturbed systems. This resonance can be harnessed for quantum enhanced sensing Lang et al. (2015); Wang et al. (2022b); Jiang et al. (2022); Pang and Jordan (2017); Zhou et al. (2020); Mishra and Bayat (2021, 2022); Yu et al. (2022); Ivanov (2021). This resonance phenomena has also been observed in gaped many-body quantum systems when the frequency of external driving matches with the lowest energy gap of the unperturbed Hamiltonian. Russomanno et al. (2016); Mishra and Bayat (2021, 2022). Various important physical features of systems with time-periodic influence can be extracted by employing Floquet formalism. For a time-periodic Hamiltonian, $H(t+\tau)=H(t)$, with periodicity $\tau=2\pi/\omega$, the solution of the Schrödinger equation follows from the Floquet theorem. The Floquet theorem gives an ansatz of the form $|\Psi(t)\rangle=\sum_{\ell}e^{-i\mu_{\ell}t}|\Phi_{\ell}(t)\rangle$. Here $\mu_{\ell}^{\prime}s$ are eigenphase (the Floquet phase), $|\Phi_{\ell}(t)\rangle$ the Floquet modes, and $\ell\in\mathbb{Z}$. Substituting the ansatz to the Schrödinger equation, $i\frac{\partial|\Psi(t)\rangle}{\partial t}=H(t)|\Psi(t)\rangle$, the Floquet modes satisfy
@@
Properties of Floquet modes $\{|\Phi_{\ell}(t)\}$ and Floquet phases $\{\mu_{\ell}\}$ have been explored for sensing periodic magnetic fields due to clusters of few qubits Lang et al. (2015). Here, the effective magnetic field due to the cluster of spins is detected using a single qubit as a sensor. The initial state of the system is $|\psi(0)\rangle{=}\frac{1}{\sqrt{2}}(\ket{\uparrow}{+}\ket{\downarrow})\otimes|\mathcal{B}(0)\rangle$, where |$\ket{\mathcal{B}(0)}$ is the initial state of the detected spin cluster at time $t=0$ and $\ket{\uparrow}$ , $\ket{\downarrow}$ are the computational basis states of the electronic spin, respectively. At later times, due to the interaction with the sensor (the qubit state), the spin cluster state becomes entangled with the sensor. The magnetic field signal is detected from the data of temporal coherence of the time-evolved state of the detected spin cluster. The coherence can be expressed in terms of the eigenvalues and eigenstates of the one-period unitary evolution operator or Floquet operator. The coherence displays a dip which occurs at the avoided crossings of the Floquet eigenstates where $e^{i\mu_{\ell}}\approx e^{-i\mu_{k}}$ or when the Floquet gap, $\Delta_{F}=\mu_{\ell}-\mu_{k}$, vanishes. Thus, by measuring the location of the coherence dip, the energy-difference between Floquet states can be observed with precision higher than the one predicted by the average Hamiltonian theory. However, this method can only be applicable for a sensing around the resonance frequencies. Sensing of arbitrary frequency fields has been proposed using a single qubit as a sensor in Floquet systems Wang et al. (2022b). In addition, periodically driven systems and the transition of Floquet eigenphases have been used for magnetic-field signal amplification Jiang et al. (2022). In Ref. Pang and Jordan (2017), an optimal control method is developed to estimate the amplitude, $B$, of a time-periodic magnetic field $B(t)=B(\cos(\omega t)\sigma_{x}+\sin(\omega t)\sigma_{z})$. In Ref. Yang et al. (2022b), a systematic variational approach for establishing optimal control was set up, where it was shown that for magnetometry, restricted controls in the form of one and two-body interactions are already sufficient to approximately achieve Heisenberg scaling. It is shown here that the QFI with respect to $B$ can achieve super-Heisenberg scaling with interrogation time, $t$, by applying an optimal control pulse as compared to the one without control. A solid state sensor with an ensemble of high density Nitrogen Vacancy (NV) centers in diamonds has been investigated both theoretically and experimentally by Floquet formalism  Zhou et al. (2020). The system, here, is a dense ensemble of electronic spins. Each electronic spin act as an independent probe for the external signal. Thus, one would expect that, by increasing the number of particles in the system, the sensitivity can be improved. The improvement, however, is severely hindered by interaction between the spins, since a large spins are confined in a small volume, the spin-spin interaction is unavoidable. By applying a series of dynamical decoupling pulses, the interaction between the spins has been suppressed, which results in a larger coherence time. In this way the standard quantum limit is shown to be restored which is lost due to spin-spin interaction.
@@ XI.1 Profiting from dissipation
Since non-classical states are crucial for countless applications and theoretical advances, it would be highly desirable to have a mechanism where instead of avoiding dissipation, one could harness the very process of dissipation to one’s advantage. The seminal work by Alfred Kastler in 1950 pointed in that direction Kastler, Alfred (1950). The process, known as optical pumping, prepares any initial quantum state into a well-defined target state via selection rules of spontaneous emission. Building on Kastler’s work, coherent population trapping was experimentally demonstrated in the 1970s by Arimondo and Orriols Alzetta et al. (1976); Arimondo and Orriols (1976). In a nutshell, a three-level energy configuration with transitions levels $|a\rangle\longleftrightarrow|b\rangle$ and $|a\rangle\longleftrightarrow|c\rangle$ and Hamiltonian $H=\hbar\Omega(|a\rangle\langle b|-|a\rangle\langle c|+h.c)$ allows for a quantum state with a zero eigenvalue (a dark state). Remarkably, such a dark state, denoted as $|\psi_{d}\rangle\sim|b\rangle+|c\rangle$, decouples from the external pumping field $\Omega$, thus allowing excitations that reach that energy level to remain there. In contrast, the bright state, denoted as $|\psi_{b}\rangle\sim|b\rangle-|c\rangle$, oscillates with the driving field $\Omega$ and the state $|a\rangle$. Indeed, this idea lies at the heart of laser cooling Aspect et al. (1988) as the ground state represents another specific quantum state. The paradigm shift of harnessing dissipation as a resource rather than a drawback has led to systematic studies of decoherence itself Myatt et al. (2000), the ability to prepare a quantum state from any arbitrary initial state Poyatos et al. (1996), the use of atomic engineered reservoirs to create two-mode Einstein-Podolsky-Rosen-entangled states Pielawa et al. (2007, 2010), protection against noise through the engineering of pointer states Carvalho et al. (2001), dissipative quantum computation Verstraete et al. (2009), the creation of many-body states and non-equilibrium quantum phases Diehl et al. (2008), thermometry of cold Fermi gases with dephasing-impurities Mitchison et al. (2020); Bouton et al. (2020); Brattegard and Mitchison (2024), quantum memories using dissipative mechanisms Pastawski et al. (2011), the generation of Bell states on an open-system quantum simulator Barreiro et al. (2011), and the possibility to enhance sensing precision nearly to the Heisenberg limit Goldstein et al. (2011).
@@ XI.3 Dissipative phase transition as a sensing resource
Intense theoretical work has been pursued to understand the critical behavior at the dissipative phase transition Rota et al. (2017); Vicentini et al. (2018); Letscher et al. (2017); Overbeck et al. (2017); Capriotti et al. (2005); Eisert and Prosen (2010); Sieberer et al. (2014); Garbe et al. (2020); Ivanov (2020), including photonic systems Carmichael (2015); Benito et al. (2016); Mendoza-Arenas et al. (2016); Casteels et al. (2016); Bartolo et al. (2016); Casteels et al. (2017); Casteels and Ciuti (2017); Foss-Feig et al. (2017); Biondi et al. (2017); Biella et al. (2017); Savona (2017); Sánchez Muñoz et al. (2018), lossy polariton condensates Sieberer et al. (2013, 2014), and spin systems Kessler et al. (2012); Lee et al. (2013); Jin et al. (2016); Lee et al. (2011); Chan et al. (2015); Maghrebi and Gorshkov (2016); Overbeck et al. (2017). In addition, fundamental limitations on criticality and driven-dissipative phase transitions in quadratic open systems, namely bilinear Hamiltonian, linear Lindblad operators, and bilinear Hermitian Lindblad operators, have been recently studied Zhang and Barthel (2022). Example of the above have been investigated for driven-dissipative Bose-Einstein condensate Zhang and Barthel (2024). Exact steady-state expressions Lawande et al. (1981); Puri et al. (1980); Bartolo et al. (2016) and approximate methods for their evaluation based on variational approaches have been pursued Weimer (2015); Raghunandan et al. (2018). The presence of phase symmetry breaking in a single qubit-laser system Fernández-Lorenzo and Porras (2017), semiclassical first-order dissipative phase transition in Kerr parametric oscillators with parity symmetry breaking Heugel et al. (2019), and second-order transitions in the two-photon Kerr resonator Di Candia et al. (2023). The universality class of driven-dissipative systems has also been extensively addressed Marcuzzi et al. (2014); Carollo et al. (2019b); Sieberer et al. (2013); Marino and Diehl (2016); Maghrebi and Gorshkov (2016) with in-depth analysis of the phase diagram and critical exponents of a dissipative transverse Ising spin chain Werner et al. (2005). From a geometric perspective, the Uhlmann curvature has been proposed to shed light upon the nature of criticality Carollo et al. (2018b), and a recent method based on the coherent anomaly approach has been proposed for determining critical exponents Jin et al. (2021).
@@
Quantum-enhanced sensitivity using the BTC probe of Eq. (76) can be evidenced in Fig. 15(a) via QFI analysis Montenegro et al. (2023). As seen from the figure, the maximum of the QFI $\mathcal{F}_{Q}^{\mathrm{max}}$ grows polynomially with the system size $N$. A fitting function of the form $\mathcal{F}_{Q}^{\mathrm{max}}=aN^{b}+c$ ($c\rightarrow 0$) reveals a coefficient $b=1.345$. Interestingly, in Fig. 15(a), a simple measurement basis of the total magnetization also scales super-linearly with the system size $N$. A fitting function of the same form as above exhibits quantum-enhanced sensitivity with a coefficient $b=1.338$. To show the performance between the QFI and the CFI, in Fig. 15(b), the ratio between their maximum values is plotted, showing an asymptotic performance of 90% —a fair fraction achieved with a sub-optimal measurement basis.
@@ XIII Non-Equilibrium Quantum Sensing: Sequential Measurements Metrology
Measurement represents another distinct feature that sets apart the classical from the quantum world. Indeed, the concept of wave-function collapse due to projective measurements on the quantum system has sparked several discussions within quantum theory itself Machida and Namiki (1980a, b); Furry (1936); Zeh (1970). In quantum many-body probes with partial accessibility, measurements can only be conducted locally on a subsystem. However, in these scenarios, even though the measurement is local, the wave function collapse is global and affects the entire system’s wave function. This intriguing phenomenon has been the subject of intensive studies Busch et al. (1990); Schmidt and Gemmer (2020); Ban (2021); Skinner et al. (2019b); Block et al. (2022a); Benoist et al. (2023); Haapasalo et al. (2016); Benoist et al. (2019); Burgarth et al. (2014); Rybár and Ziman (2015), such as a unique feature for quenching many-body systems Bayat (2017); Pouyandeh et al. (2014); Ma et al. (2018); Bayat et al. (2018), leading to potential new types of phase transitions Choi et al. (2020); Feng et al. (2023); Minato et al. (2022); Skinner et al. (2019a); Gullans and Huse (2020); Block et al. (2022b). These measurement-induced phase transitions have also been recently linked with quantum enhanced sensitivity (Paviglianiti and Silva, 2023; Di Fresco et al., 2024).
@@
where $\bm{\sigma}_{j}{=}(\sigma_{j}^{x}{,}\sigma_{j}^{y}{,}\sigma_{j}^{z})$ represents a vector composed of Pauli matrices acting on qubit site $j$, $J$ denotes the exchange interaction, and $\bm{B}{=}(B_{x}{,}0{,}B_{z})$ stands for the local magnetic field to be estimated located at first site. For the sake of simplicity, a ferromagnetic state $|\psi(0)\rangle=|\downarrow\downarrow\downarrow\ldots\rangle$ is probed at site $N$ in the $\sigma_{z}$ basis at regular time intervals $J\tau=N$. The latter allows enough time for the quantum correlations to spread throughout the spin chain. It has been shown Montenegro et al. (2022) that a synchronization-like magnetization dynamics occurs between the first and last sites. This opens up the possibility of extracting information about the local magnetic field (located at the first site) by observing the dynamics of the readout qubit (located at the last $N$ site). Remote sensing holds special significance in biology, where non-invasive sampling of organic tissues is desirable to protect the biological material (see quantum illumination techniques Albarelli et al. (2020); Casariego et al. (2022)).
@@
where $\mathcal{F}_{\lambda}^{(j)}$ represents the CFI at step $j$, $\Delta\mathcal{F}_{\lambda}^{(n)}$ is the increment of the CFI after performing one more measurement following the recording of $(n{-}1)$ measurements, and $f_{\lambda}^{\boldsymbol{\gamma},(n)}$ is the CFI obtained from the $n$-th measurement $p(\gamma_{n})$ in trajectory $\boldsymbol{\gamma}$. Utilizing this approach, one can examine the behavior of the CFI for arbitrarily large numbers of sequential measurements. Conducting local projections on a quantum many-body probe leads to probe’s finite memory, attributed to repeated local projections and free evolution, resulting in a rank-1 matrix Yang et al. (2023). This probe’s finite memory is linked to the transition from super-linear to linear scaling of CFI. After losing memory of an early state, CFI scales linearly with $n_{\mathrm{seq}}$, limiting the extractable information capacity of the quantum probe in sequential measurements.
For the quantum many-body probe described in Eq. (82), Figs. 18(a)-(b) demonstrate a correspondence between the transition from super-linear to linear scaling of CFI and the probe’s finite memory. Specifically, the CFI increment $\Delta F_{B}^{(n_{\mathrm{seq}})}$ becomes nearly constant at around the same $n_{\mathrm{seq}}$ where the averaged fidelity $\langle F\rangle_{\mathrm{traj}}$ also becomes nearly constant. This averaged fidelity measures the similarity between two distinct initial states following the same trajectory Yang et al. (2023) (a single trajectory is composed of $n_{\mathrm{seq}}$ measurement steps). In Fig. 18(c), the CFI $\mathcal{F}_{B}^{(n_{\mathrm{seq}})}$ is plotted as a function of $n_{\mathrm{seq}}$. This plot clearly shows a transition from non-linear to linear behavior, in agreement with the quantum-enhanced sensing shown for short sequential measurements Montenegro et al. (2022) and the linear scaling with indirect methods for large $n_{\mathrm{seq}}$ Burgarth et al. (2014). The Monte Carlo-based methodology discussed above for arbitrarily large $n_{\mathrm{seq}}$ it has been addressed for both quantum many-body probes and light-matter probes Yang et al. (2023).
Figure 18: Sequential measurements metrology for the quantum-many body probe of Eq. (82) and large number of consecutive measurements $n_{\mathrm{seq}}$: $(a)$ CFI increment $\Delta F_{B}^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$, $(b)$ Fidelity averaged over $10^{4}$ different trajectories (each trajectory evaluates the fidelity between two distinct initial states following the same trajectory) as a function of $n_{\mathrm{seq}}$, $(c)$ CFI $\mathcal{F}_{B}^{(n_{\mathrm{seq}})}$ as a function of $n_{\mathrm{seq}}$. Figure taken from Yang et al. (2023).